OpenStudy (anonymous):

2cos^2x+sinx-2=0 Solve the equation on the interval [0,2pi]

OpenStudy (anonymous):

@Directrix

OpenStudy (kc_kennylau):

Write cos^2x in (1-sin^2x)

OpenStudy (anonymous):

What next ?

OpenStudy (kc_kennylau):

Let S = sinx

OpenStudy (anonymous):

What does that mean?

OpenStudy (kc_kennylau):

Well, firstly we have \(2(1-\sin^2x)+\sin x-2=0\)

OpenStudy (kc_kennylau):

Now we have \(2(1-S^2)+S-2=0\)

OpenStudy (kc_kennylau):

Do you understand?

OpenStudy (anonymous):

Yes

OpenStudy (kc_kennylau):

Then solve the quadratic equation

OpenStudy (anonymous):

Disturbute 2?

OpenStudy (kc_kennylau):

yep

OpenStudy (anonymous):

I stopped at \[2\sin(x)^2+\sin(x)\]

OpenStudy (kc_kennylau):

Well I thought it's an equation

OpenStudy (anonymous):

Meaning?

OpenStudy (kc_kennylau):

where's the equal sign

OpenStudy (anonymous):

2sin(x)^2+sin(x)=0

OpenStudy (kc_kennylau):

it should be - instead of +

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Oh ok

OpenStudy (kc_kennylau):

then solve that equation

OpenStudy (kc_kennylau):

You have \(S^2-2S=0\) right?

OpenStudy (anonymous):

yes

OpenStudy (kc_kennylau):

then solve this quadratic equation

OpenStudy (anonymous):

Factor?

OpenStudy (kc_kennylau):

yep

OpenStudy (anonymous):

\[sinx(1-2sinx)=0\]

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

Sinx=0

OpenStudy (anonymous):

Sinx=-1/2

OpenStudy (kc_kennylau):

yep

OpenStudy (kc_kennylau):

well, not -1/2 but +1/2

OpenStudy (anonymous):

It'll be negative

OpenStudy (kc_kennylau):

why?

OpenStudy (anonymous):

1-2sinx=o

OpenStudy (anonymous):

That's how

OpenStudy (kc_kennylau):

well, 1-2sinx=0, 1=2sinx, 1/2=sinx

OpenStudy (anonymous):

When you the put the one to the right it becomes negative

OpenStudy (kc_kennylau):

1-2sinx=0 1-2sinx+2sinx = 0+2sinx 1=2sinx

OpenStudy (anonymous):

What would be the x

OpenStudy (kc_kennylau):

well, sinx=0 or sinx=1/2

OpenStudy (kc_kennylau):

what can x be?

OpenStudy (anonymous):

Pi, pi/6

OpenStudy (kc_kennylau):

there should be 5 answers

OpenStudy (kc_kennylau):

they're 0,pi,2pi and pi/6,5pi/6

OpenStudy (kc_kennylau):

i got to go now